# EIKE Workshop at JOWO

### v1

@@ Abstract
Metaphorical meaning is not a flat mapping between concepts, but a complex cognitive phenomenon that integrates multiple levels of interpretation. In this paper, we propose a stratified model of metaphor processing that treats meaning as an onion: a multi-layered structure comprising (1) content analysis, (2) conceptual blending, and (3) pragmatic intentionality. This three-dimensional framework allows for a richer and more cognitively grounded approach to metaphor interpretation in computational systems. At the first level, metaphors are annotated through basic conceptual elements. At the second level, we model conceptual combinations, linking components to emergent meanings. Finally, at the third level, we introduce a pragmatic vocabulary to capture speaker intent, communicative function, and contextual effects, aligning metaphor understanding with pragmatic theories. By unifying these layers into a single formal framework, our model lays the groundwork for computational methods capable of representing metaphorical meaning beyond surface associations—toward deeper, more context-sensitive reasoning.
@@ 1 Introduction
    Theoretical fragmentation. Competing accounts (e.g., CMT vs. interactional or embodiment theories) illuminate different aspects of metaphorical phenomena but rarely integrate pragmatics—what a metaphor does in conversation—often goes unmodeled despite Speech Act Theory being a long-standing account of these processes and has been widely adopted in the literature [4].
@@
The starting point for this framework lies in the observation that implicit linguistic meaning operates on a different level from that of surface-level, literal meaning, as exemplified in Figure 1. Conceptual and pragmatic meanings are not directly encoded in the literal expression; rather, they are inferred through processes of contextualization and implicature.
Figure 1: The layer of literal linguistic expression vs. the implicit knowledge one.
@@
When we consider, for example, the sentence we are wasting our time, the conceptual interpretation—where notions such as finite resources, waste, and time emerge, and where an analogical operation groups features of these concepts—is not directly expressed by the literal sentence. Similarly, the utterance’s potential pragmatic function—i.e., the way that the utterance is intended—is also not explicitly encoded. This distinction reflects our view of meaning as a multi-level phenomenon, where literal meaning and interpretation can be separated from conceptual and pragmatic ones, which remains implicit. Although stratified or hierarchically structured models of meaning are well established in the literature on Pragmatics (see, for example, Grice [7]), computational metaphor analysis and pragmatics have not yet developed a systematic bond, despite the fact that such a connection could greatly benefit the study of these matters.
@@
Therefore, from this theoretical perspective we propose a framework focused on the implicit level of meaning that aims to study metaphors as cognitive tool in context, which does not exist as an abstract operation, separate from its usage or from concrete linguistic experience. A simultaneous cognitive and pragmatic analysis would in fact support the view of metaphor as a figurative language element that effectively works—that is, it achieves communicative efficacy—when it is perceived as conceptually appropriate and functionally aligned with the speaker’s pragmatic goal. In this view, the implicit knowledge conveyed by metaphors can be processed more efficiently in computational systems if conceived as a multi-layered entity, like an onion. Each of our onion layer can be ontologically represented, providing a structured and functional approach to modeling metaphorical knowledge that aligns with our analytical goals. If the outermost layer correspond to the level of basic conceptual mapping, moving inward reveals the level at which the analogical conceptual operations take place, and deeper still, the pragmatic intention that motivated the entire utterance. By connecting CMT’s cognitive account of metaphor with Speech Act Theory–inspired pragmatics, our stratified framework aim to capture both what metaphors map and what they accomplish in communication. This unified model lays the groundwork for richer datasets, standardized evaluation, and computational systems that better mirror how humans unlock implicit knowledge through metaphor.
@@
ello this is a trial text
1.1 Sottosezione
1.1.1 Sottosottosezione
paragrafo=terza sottosezione
ciaoooooooooooooooo
testo testo testo codice maiuet questo va aggiunto hhhhhokkkkhh ooooooo
  • non numerato
  1. 1.
    numerato
As this paper said, [25]
“virgolette”
Notanota111nota a pié di pagina
citazione lunga
@@ 2 Related Work
In this section, we describe the background of our proposal and discuss the challenges related to existing works on metaphor processing.
@@ 2.1 Metaphor theories
Conceptual Metaphor Theory (CMT), developed by G. Lakoff and M. Johnson, posits that metaphors map a source domain onto a target domain via systematic correspondences, enabling abstract reasoning through familiar experiential structures [2]. Conceptual Blending Theory (CBT) extends this view by introducing a generic “blend” space that selectively inherits elements from both input domains according to a blending criterion or key property, such as yellow when we say “golden hair”, evoking that golden is yellow and shiny like the hair [8]. CBT is regarded as a valid computational approach also by the Categorization theory [9]. This account sees metaphors as category statements where the source acquires a categorical meaning, more abstract than its literal meaning. For example, “golden” in “golden hair” denotes the category of “shiny, yellow things”. In this view, conceptual blending can be used to extract the abstract meaning of the source and combine it with the meaning of the target. These frameworks emphasize that metaphor comprehension relies on shared background knowledge (or frames), which Fillmore’s frame semantics formalizes by associating lexical items with structured role–filler expectations [10]. The complementary relationship between CMT, CBT, and frame semantics highlights that metaphorical meaning emerges not merely from lexical similarity but from dynamic frame activation and role alignment within a community’s commonsense knowledge [4]. Beyond commonsense or prototypical knowledge, recent theories of metaphors have noted the lack of inclusion of personal and sometimes contextual aspects that influence knowledge acquisition and interchange. In fact, the experiential dimension of metaphor has traditionally been downplayed, with research focusing primarily on metaphors as a mental and individual achievement. Researchers have so far paid little attention to context and the collaborative production of metaphoric language [11].
@@ 2.5 The pragmatics perspective
Pragmatics, as the study of the meaning of linguistic signs in context—that is, in their actual use—deals primarily with implicit linguistic knowledge. The type of meaning it investigates, the pragmatic message, is not encoded in any direct way in the literal utterance. One can say something that literally means one thing while actually intending something entirely different. For instance, saying to someone on the subway, “You’re standing on my foot”, means likely not wanting to describe the situation to them, but rather asking them to move [37]. The utterance may contain only hints as to how the pragmatic meaning should be interpreted, such as tone of voice, and it depends on a variety of extra-textual and extra-linguistic elements, including context, linguistic conventions, and socio-cultural norms. These pragmatic cues, along with contextual dependence, must be correctly interpreted, something that requires what is known as pragmatic competence. Utterances that contain metaphors can, of course, be described in pragmatic terms, and can further be understood as speech acts. Speech Act Theory (SAT) [38] [39] considers language and linguistic utterances not merely as expressions of mental operations, like articulations of thoughts, but as actions in themselves. As actions they produce effects in the world, while pursuing their speaker’s intentions. These are acts such as describing, ordering, pleading, mocking, or, through conventional formulas, marrying two people or sentencing someone.
Computational pragmatics deals mainly on how types of speech acts, conceived as categories of illocutionary acts—that is, acts defined by the speaker’s intention, by how they mean the literal sentence they utter—are assigned to utterances, framing this as a problem of context dependence. This is far from a trivial task: there is no deterministic relationship between clause types and illocutionary force, in the sense that imperative clauses are always commands, interrogative clauses always queries, and declarative clauses always assertions. Depending on factors like the power dynamics between speaker and addressee, and their communicative goals, an utterance may emerge as a complex blend of request, suggestion, and command. A “metaphorical speech act” would presumably involve a similar blend of illocutionary forces. According to certain frameworks, such as those proposed by Popa-Wyatt, one can distinguish primary and secondary illocutionary acts, sometimes nested within each other, as, for instance, in the case of an ironic act embedded inside a metaphorical one [40]. Though, speech act labeling should not be valuable as an exercise in itself, but only insofar as it helps identify the communicative purposes of an interaction, and, likewise, categorization of illocutionary acts should be interpreted functionally, rather than as fixed descriptive categories. Other main challenges in computational pragmatics include identifying and tracking the effects that utterances have on context, modeling context itself, and formalizing cultural conventions and intentionality, all of which are crucial for a pragmatic study of metaphor as well. In truth, in computational pragmatics, the pragmatic understanding of metaphor—why it is used, in what context, and with what effect—remains only partially addressed, despite the advances of contextual neural models in recent years. As we have seen, metaphor is often treated within the realm of automatic recognition and pragmatic-inferential interpretive approaches in the tradition of CMT, which aim to capture conceptual mapping. However, to be truly robust, these approaches require a clear representation of communicative goals. There are also attempts to address metaphor from the perspective of intention and dialogue, such as models of Theory of Mind based on beliefs and intentions, sometimes inspired by the Relevance Theory (RT), and adapted computationally [41]. In this view, metaphor is treated as a pragmatic tool conveying implicatures and implicit evaluations. Relevance Theory offers several strengths in capturing implicit meaning, particularly in its emphasis on inferential comprehension and the cognitive principle that human communication aims at maximal relevance with minimal processing effort. As argued by Gibbs [42], RT provides a flexible framework for understanding metaphor as an expression of speaker intention that relies on the hearer’s ability to infer non-literal meanings from contextual cues and assumptions. Unlike Speech Act Theory, which classifies utterances according to fixed illocutionary forces, RT focuses on the dynamic interplay between context, cognition, and communicative intention. While SAT offers a useful structural categorization of communicative acts, RT excels in modeling how listeners derive meaning beyond what is explicitly said. A fruitful direction for research lies in integrating these two perspectives: leveraging the inferential depth of Relevance Theory to handle non-literal and context-sensitive meaning, while also drawing on the intentional structure of Speech Act Theory to account for the performative dimension of metaphorical utterances. This is precisely what we aim to pursue, in order to better model metaphor as a communicative act endowed with pragmatic force.
@@ 3 How meaning is an onion: a proposal of three-dimensional meaning representation
In this section, we explain the main thesis of our research proposal and the three-dimensional meaning framework. It serves as a vocabulary that, upon discussion in the community, can lay the ground for future ontological implementations.
3.1 Metaphors as speech acts
Studying metaphors as speech acts can be particularly productive for at least two reasons. First, a pragmatic interpretation of metaphors can contribute significantly to tasks such as automatic language processing and interpretation, by making explicit the non-literal elements that are crucial in multimodal contexts, for example, the context of the speech act itself and the pragmatic cues, whether textual (e.g., punctuation) or somatic (e.g., tone of voice, physical gestures). It also facilitates the study of the speaker’s intention, as inferred through implicit pragmatic reasoning. Importantly, intention is treated as something that emerges from the analysis itself. From a computational perspective, this means modeling intention as an inferred property of communication, rather than a presupposed internal condition. Second, if metaphor-containing utterances are speech acts, then we can also examine the pragmatic effects of metaphor as a figure-of-speech device. As part of linguistic style, rhetorical figures like metaphor modulate the perception of speaker intention, and therefore influence the persuasive or expressive force of the utterance. Shifts in stylistic tone can dramatically impact the pragmatic force of a speech act: a metaphor used within a given illocutionary intention can either reinforce or attenuate that intention. A command may sound more polite, a request more engaging, or a threat more forceful. This makes metaphor a particularly interesting case for the study of pragmatic implicit meaning, precisely because it enables the perception of linguistic manipulation through conceptual evocation, which directly relates to persuasion, that is, pragmatic effectiveness—how well the utterance serves the speaker’s communicative goals and fits the interlocutor’s sensitivity and context.
To the analysis of illocutionary intention, one can add the analysis of the perlocutionary act, that is, the name for the speech act performed by saying something. It concerns the effect an utterance or image has on the listener (e.g., convincing, frightening, provoking thought). This level of analysis goes beyond the semantic-pragmatic interpretation of an expression and the linguistic competence involved, merging with psychological factors. However, understanding it means evaluating not only what a metaphor means, but also what a metaphorical utterance does—in line with the treatment of utterances as linguistic actions, and as rhetorical figures, metaphors can aim at linguistic persuasion, including the evocation of emotional responses.
@@ 3.2 The layered perspective
Understanding the meaning of an expression requires a multi-level analysis of its content, of the conceptual effects it produces and of the pragmatic context in which it is produced. We propose a three-layered approach to meaning representation that allows to perform such analyses in parallel.
@@
3.2.1 Layer 1: The conceptual content layer
The conceptual content layer includes metadata of the object we are analyzing: domain, provenance, connections to frames, etc. It also includes annotator metadata.
@@ 3.2.2 Layer 2: The conceptual combination layer
The context of this layer is given at least by the annotation of source and target domains and is fully represented according to the Blending Ontology (see Section 2) with the following elements:
  • Blendable (source and target domains)
  • Blending (the blending principle)
  • Blended (the actual blend)
@@ 3.2.3 Layer 3: The pragmatic layer
This third layer is so far unexplored with respect to the first two and defines the pragmatic context of an utterance. This approach extends across modalities: beyond text and images, it can, for example, recognize tone of voice in speech. Specifically, we propose the following categories to take into account:
@@
  • Illocutionary Act: As discussed above, this refers to the act that expresses how speakers intend their literal utterance to be understood—thus representing the key element from which speaker intention can be inferred. In our framework intention is not treated as an independent mental state, but rather as something that emerges from the function and attitude of the illocutionary act. In other words, the act itself manifests the communicative force and intent in the context of interaction, without the need to presuppose an internal mental representation. Following widely adopted classifications—such as Searle’s taxonomy—illocutionary acts include assertive, directive, commissive, expressive, and declarative types, but as we have seen, broader approaches may also consider metaphorical or ironic acts as legitimate illocutionary categories. Given that communicative acts often involve a composite of multiple illocutionary forces, our research focuses specifically on elements that can be traced back to directive illocutionary acts.
@@
  • Perlocutionary Act: The speech act concerning the psychological or emotional effects an utterance or image has on the listener/interpreter. This includes both intended and unintended responses. As rhetorical devices, metaphors can function persuasively by evoking emotional reactions and shaping the audience’s response beyond the literal content.
    • Perlocutionary Effect: The emotions evoked by the metaphor in the interpreter, categorized according to the Emotion Frame Ontology [45]. These emotional responses may vary and are optional: in some cases a metaphor may evoke no particular emotion.
    • Efficacy: A measure (on a Likert scale from 1 to 5) of the metaphor’s effectiveness, namely its capacity for persuasion relative to the presumed illocutionary intention, its appropriateness to the perceived context, and the interpreter’s personal sensitivity.
@@
    • Pragmatic/Visual Clues: Indicators that suggest the speaker’s or communicator’s attitude toward the metaphor’s referent.
@@
As an example, consider again the image shown in Figure 6. The speaker’s evaluative stance is unmistakably negative toward the analogy’s object: smoking is framed as a lethal threat and self-destructive act. This negative attitude is conveyed both conceptually (via the metaphor) and visually (via design choices). The utterance, although visual, functions as a directive illocutionary act because it is embedded in the context of anti-smoking public campaigns and visually represents a cigarette as a bullet. This metaphor implies a communicative function—discouraging smoking behavior by equating it with an act of self-inflicted violence. From this function, the speaker’s intention can be inferred as emerging from the pragmatic force and evaluative stance of the act itself, as to prevent smoking behavior by highlighting its deadly consequences through the conceptual mapping of smoking is shooting oneself. As for the directive kind, the act falls on the highly assertive end of the directive spectrum, operating with a prohibitive force. It essentially issues a prohibition, commanding viewers not to smoke. This is achieved not through explicit verbal instruction but via the emotional and conceptual weight of the metaphorical image. The perlocutionary effects include fear, concern, and shock as viewers process the visual equation of cigarettes with ammunition. The metaphor’s efficacy is potentially high due to the stark, unambiguous nature of the weapon imagery and its cultural associations with death and violence. Visual clues reinforce the negative attitude through the absence of color, and the conceptual mapping between shooting a gun and smoking a cigarette is underlined by the presence of smoke. The overall visual tone of voice is dramatic and alarming, designed to interrupt habitual thinking about smoking through visceral impact rather than rational argument. This example of analysis is outlined in Figure 5.

### v2

# CEUR requirements: “Copyright © 2025 for this paper by its authors. Use permitted under Creative Commons License Attribution 4.0 International (CC BY 4.0)5

@@ Abstract
Metaphorical meaning is not a flat mapping between concepts, but a complex cognitive phenomenon that integrates multiple levels of interpretation. In this paper, we propose a stratified model of metaphor processing that treats meaning as an onion: a multi-layered structure comprising (1) contextual information, (2) conceptual blending analysis, and (3) pragmatic analysis. This three-dimensional framework allows for a richer and more cognitively grounded approach to metaphor interpretation in computational systems. At the first level, metaphors are annotated through contextual metadata. At the second level, we model conceptual combinations, linking components to emergent meanings. Finally, at the third level, we introduce a pragmatic vocabulary to capture speaker intent, communicative function, and contextual effects, aligning metaphor understanding with pragmatic theories. By unifying these layers into a single formal framework, our model lays the groundwork for computational methods capable of representing metaphorical meaning beyond surface associations—toward deeper, more context-sensitive reasoning.
@@ 1 Introduction
    Theoretical fragmentation. Competing accounts (e.g., CMT vs. interactional or embodiment theories) illuminate different aspects of metaphorical phenomena but rarely integrate pragmatics—what a metaphor does in conversation—often goes unmodeled despite Speech Act Theory (SAT) being a long-standing account of these processes and widely adopted in the literature [4].
@@
The starting point for this framework lies in the observation that, for analytical purposes, implicit linguistic meaning can be treated as operating on a different level from that of explicitly encoded, surface-level meaning. Conceptual and pragmatic meanings are not directly encoded in the expression itself; rather, they are inferred through processes of contextualization and implicature. Consider, for example, the sentence we are wasting our time: the conceptual interpretation—where notions such as finite resources, waste, and time emerge, and where an analogical operation groups features of these concepts—is not directly expressed by the literal sentence. Similarly, the utterance’s potential pragmatic function—i.e., the way that the utterance is intended—is also not explicitly encoded. This distinction reflects a view of meaning as a multi-level phenomenon—not necessarily intended as a model of how language inherently works, but as a potentially effective way of structuring meaning for computational processing, where explicitly encoded meaning and interpretation can be distinguished from conceptual and pragmatic ones, which remain implicit. Although stratified or hierarchically structured models of meaning are well established in the literature on Pragmatics (e.g. in Grice [7]), computational metaphor analysis and pragmatics have yet to establish a systematic connection—a link that could substantially advance research in both domains.
@@
Therefore, from this theoretical perspective we propose a framework focused on the implicit level of meaning that aims to study metaphors as cognitive tools in context, which do not exist as an abstract operation, separate from its usage or from concrete linguistic experience. A simultaneous cognitive and pragmatic analysis would in fact support the view of metaphor as a tool that effectively works—that is, it achieves communicative efficacy—when it is perceived as conceptually appropriate and functionally aligned with the speaker’s pragmatic goal, in addition to serving as models that reflect the cognitive structuring of experience, shaping both reasoning and action [1]. In this view, the implicit knowledge conveyed by metaphors can be processed more fruitfully in computational systems if conceived as a multi-layered entity, like an onion. Each of our onion layer can be ontologically represented, providing a structured and functional approach to modeling metaphorical knowledge that aligns with our analytical goals. If the outermost layer correspond to the level of contextual information mapping, moving inward reveals the level at which the analogical conceptual operations take place, and deeper still, the pragmatic intention that motivated the entire utterance. By connecting CMT’s cognitive account of metaphor with SAT–inspired pragmatics, our stratified framework aim to capture both what metaphors map and what they accomplish in communication. This unified model lays the groundwork for richer datasets, standardized evaluation, and computational systems that better mirror how humans unlock implicit knowledge through metaphor.
@@ 2 Related Work
In this section, we describe the theorical background for our proposal and discuss the challenges related to existing works on metaphor processing.
@@ 2.1 Metaphor theories
CMT, developed by Lakoff and Johnson, posits that metaphors map a source domain onto a target domain via systematic correspondences, enabling abstract reasoning through familiar experiential structures [2]. Conceptual Blending Theory (CBT) extends this view by introducing a generic “blend” space that selectively inherits elements from both input domains according to a blending criterion or key property, such as yellow when we say “golden hair”, evoking that golden is yellow and shiny like the hair [8]. CBT is regarded as a valid computational approach also by the Categorization theory [9]. This account sees metaphors as category statements where the source acquires a categorical meaning, more abstract than its literal meaning. For example, “golden” in “golden hair” denotes the category of “shiny, yellow things”. In this view, conceptual blending can be used to extract the abstract meaning of the source and combine it with the meaning of the target. These frameworks emphasize that metaphor comprehension relies on shared background knowledge (or frames), which Fillmore’s frame semantics formalizes by associating lexical items with structured role–filler expectations [10]. The complementary relationship between CMT, CBT, and frame semantics highlights that metaphorical meaning emerges not merely from lexical similarity but from dynamic frame activation and role alignment within a community’s commonsense knowledge [4]. Beyond commonsense or prototypical knowledge, recent theories of metaphors have noted the lack of inclusion of personal and sometimes contextual aspects that influence knowledge acquisition and interchange. In fact, the experiential dimension of metaphor has traditionally been downplayed, with research focusing primarily on metaphors as a mental and individual achievement. Researchers have so far paid little attention to context and the collaborative production of metaphoric language [11].
@@ 2.5 Metaphors as speech acts
Pragmatics, as the study of the meaning of linguistic signs in context—that is, in their actual use—deals primarily with implicit linguistic knowledge: the type of meaning it investigates, the pragmatic message, is not encoded in any direct way in the literal utterance. One can say something that literally means one thing while actually intending something entirely different. For instance, saying to someone on the subway, “You’re standing on my foot”, means likely not wanting to describe the situation to them, but rather asking them to move. The utterance may contain only hints as to how the pragmatic meaning should be interpreted, such as a specific tone of voice, and it depends on a variety of extra-textual and extra-linguistic elements, including context, linguistic conventions, and socio-cultural norms.
Utterances that contain metaphors can, of course, be described in pragmatic terms; however, pragmatic approaches to metaphor depend on the different approaches to the interpretation of metaphor itself. In that of two leading figures in the pragmatics literature and of SAT such as Grice and Searle [37, 38] there is the idea that the metaphorical interpretation presupposes and is derived from the literal interpretation through a “decoding” of a secondary level of meaning, as if metaphors and other figures of speech were some exceptional or supplementary use of language. The assumption of a clear-cut distinction between the literal interpretation and the metaphorical interpretation of an utterance is instead abandoned in the approach to metaphor of Sperber and Wilson [39, 40], the theorists of Relevance Theory (RT). RT, whose strength lies in its understanding of how language works through its emphasis on the inferential comprehension of non-literal meanings from contextual cues and assumptions, and on the cognitive principle that human communication aims at maximal relevance with minimal processing effort, treats metaphor as a pragmatic tool conveying implicatures and implicit evaluations.
The cognitive approach to the study of metaphors has been increasingly followed after the rise of CMT, and even in pragmatics, studies have focused their research on the many and varied effects that metaphor has on cognitive processes (e.g., [41, 42]). However, in computational pragmatics the field still appears to be very open: the pragmatic processing and understanding of metaphor—why it is used, in what context, and with what effect—remains only partially addressed, despite the advances of contextual neural models in recent years. Focusing exclusively on cognitive aspects may also obscure others, such as the performative dimension of utterances, which is instead well highlighted by SAT. Speaking of utterances as speech acts, in fact, means considering language and linguistic utterances not merely as expressions of mental operations, like articulations of thoughts, but as actions in themselves. As actions, while pursuing their speaker’s intentions, they produce effects in the world, whether intended or not, and carry out acts such as describing, ordering, pleading, or, through conventional formulas, marrying two people or sentencing someone.
A central aspect in explaining pragmatic meaning is the intention attributed to an utterance, the illocution—that is, the way in which the speaker intends the literal sentence they utter, for example, saying something with the intention of making a request or giving an order. One of the main goals of computational pragmatics inspired by SAT—namely, studying how to automatically assign an illocutionary act to the utterance intended to express it, framing this as a problem of context dependence—is complicated by several factors. First, there is the challenge of formalizing intentional, conventional, or otherwise contextual aspects that are extra-linguistic, along with all the choices that such formalization entails. Second, there is no deterministic relationship between clause types and illocutionary force: imperative clauses are not invariably commands, interrogative clauses are not always queries, and declarative clauses are not necessarily assertions. Finally, any illocutionary classification, however widely shared, will inevitably fall short of capturing the compositionality of the intentions at play in natural language. Depending on factors such as the power dynamics between speaker and addressee and their communicative goals, an utterance may emerge as a complex blend of illocutionary forces; and the same applies in the case of a “metaphorical speech act”—understood here not as a distinct illocutionary category, but as a speech act that is in some way metaphorical.
Thus, as noted by Michelli, Tong and Shutova [43], the literature on metaphor intention is fragmented: there is still a lack of a systematic and comprehensive account of intentions behind metaphor use and an operationalised framework enabling the annotation of such intentions in linguistic data. However, explaining the communicative role of metaphors in terms of taxonomies of intentions, as they propose, may risk to perpetuate a classificatory game that views intentions as isolated. Another thing to note is that, although there is not a common notion of intention shared among all metaphor scholars, intentions are typically formalized as prior intentions, that is, as representations in the speaker’s mind of their communicative goals, especially in approaches that address metaphor from the perspective of models of Theory of Mind based on beliefs and intentions. As Gibbs [44] points out, conceiving of intentions as individual mental states makes them opaque, since agents are not always aware of the causes of their behaviour; this, in turn, can lead to computational abstractions that lack real explanatory value for illocutionary intention.
Studying metaphors as speech acts, instead, means treating intention not as a mental state, but as a feature attributed to linguistic acts, in line with the philosophy of action outlined by Austin in his linguistic analyses [45, 46]. This means that intentions, rather than being the reasons that speakers may provide when asked why they resorted to certain metaphors, are instead the intentions expressed by the utterance itself (thus interpreting an utterance as a command rather than a statement). Consequently, it is not immediately necessary to presuppose mental states beyond linguistic analysis. From a computational perspective, this entails modelling intention as an inferred property of the communication rather than as a presupposed mental condition—certainly a challenge, but also a potentially valuable contribution to the field of computational pragmatics. Second, if metaphor-containing utterances are speech acts, they also produce effects on the communicative context, and part of which derive from the metaphor as a cognitive and stylistic device. The pragmatic effectiveness of a metaphorical speech act—how well the utterance serves the speaker’s communicative goals and fits the interlocutor’s sensitivity and context—is shaped by the stylistic tone conveyed through rhetorical figures and by the strength of the conceptual evocation created by the metaphor itself. Within a given illocutionary intention, a metaphor can either reinforce or attenuate that intention: a command may sound more polite, a request more engaging, or a threat more forceful; conversely, if the metaphor is perceived as inappropriate, its expressive or persuasive force may be diminished. This suggests that, beyond analysing illocutionary intention, it is also relevant to consider the perlocutionary level (to use Austin’s terminology, the perlocutionary, the name for the speech act performed by saying something) of a metaphorical speech act, which concerns the effect an utterance or image has on the listener (e.g., convincing, frightening, provoking thought). These effects can be cognitive—often more difficult to study—but also at a more immediate at the emotional-psychological level, since, as figures of speech, metaphors can aim at linguistic persuasion, including the evocation of emotional responses.
@@ 3.1 The layered perspective
We will discuss now the main thesis of our research proposal and the three-dimensional meaning framework. It serves as a vocabulary that, once discussed within the community, can lay the groundwork for future computational implementations. The main aim for the framrwork is to provide, as an overall framework, the interpretation of metaphor as both a cognitive tool and a speech act.
@@
3.1.1 Layer 1: The external context layer
The conceptual content layer includes metadata of the communicative object we are analyzing: domain, provenance, connections to existing knowledge bases, etc. It also includes annotator metadata.
@@ 3.1.2 Layer 2: The conceptual combination layer
The context of this layer is given at least by the annotation of source and target domains and is fully represented according to the Blending Ontology (see Section 2) with the following elements: (i) Blendable (source and target domains); (ii) Blending (the blending principle); (iii) Blended (the actual blend).
@@ 3.1.3 Layer 3: The pragmatic layer
Compared with the first two layers, the third remains largely under-investigated from a computational perspective. Our aim is to discuss a categorization that would allow for its effective implementation, also with a view to the composition and annotation of a dataset. The idea, in fact, is to propose a human annotation campaign on a dataset—such as one consisting of metaphors found in images, as in the example discussed so far—and then to assess the behavior of an automatic system on the same task.
As seen in Section 2.5, works on the communicative role of metaphors in a pragmatic sense tend to explain it in terms of the intentions or discourse goals they are meant to achieve, proposing taxonomies of intentions [43], that often assumes intentions as mental states. For the analysis of intention as an illocutionary act, we reuse the standard categories of SAT, focusing on defining the role of metaphorical linguistic action [^9] The general classification of illocutionary acts (or language functions) has been a recurring topic in the philosophy of language literature, but here we are not concerned with defending a particular classification. We adopt an approach that would allow us to flexibly assess the composition of forces within an utterance containing a metaphor, focusing on the example of directives—the linguistic category of requests or orders typically derived from imperative sentences and found across most languages and linguistic cultures [49]. According to certain theoretical frameworks, such as those proposed by Popa-Wyatt, one could also distinguish primary and secondary illocutionary acts, sometimes nested within each other, as, for instance, in the case of an ironic act embedded inside a metaphorical one [50], but given the preliminary nature of this operational framework, we focus on annotating only the forces perceived as primary..
As shown, a perlocutionary analysis aligns with evaluating not only what a metaphor means, but also what it does, treating utterances as actions that produce effects in the communicative context and in listeners, whether intended or not. We propose capturing these effects through a first, immediate psychological evaluation, and extend this approach across modalities, including the recognition of tone of voice in speech as a cue to pragmatic intention.
We therefore propose the following categorization:
@@
  • Illocutionary Act: The act that expresses how speakers intend their literal utterance to be understood—thus representing the key element from which utterance’s intention can be inferred. Following classifications such as Searle’s taxonomy, illocutionary acts can include assertive, directive, commissive, expressive, and declarative types. Communicative acts often involve a composite of multiple illocutionary forces, but our case study focuses specifically on elements that can be traced back to directive illocutionary acts as the principal pragmatic force perceived by annotators.
@@
  • Perlocutionary Act: The act concerning the effects an utterance or image has on the listener/annotator, here analyzed as psychological or emotional. This includes both intended and unintended responses, which we propose to capture as annotators’ evaluations.
    • Perlocutionary Effect: The emotions evoked by the metaphor in the annotator, categorized according to the Emotion Frame Ontology [51]. These emotional responses may vary and are optional: in some cases a metaphor may evoke no particular emotion.
    • Efficacy: A measure (on a Likert scale from 1 to 5) of the metaphor’s effectiveness, namely its capacity for persuasion relative to the presumed illocutionary intention, its appropriateness to the perceived context, and the annotator’s personal sensitivity.
@@
    • Pragmatic/Visual Clues: Indicators that suggest the speaker’s or communicator’s attitude toward the metaphor’s referent, e.g. use of specific colors.
@@
As an example, consider again the image shown in Figure 5. The speaker’s evaluative stance can easily be perceived by an annotator as negative toward the analogy’s object: smoking is framed as a lethal threat and self-destructive act. This negative attitude is conveyed both conceptually (via the metaphor) and visually (via design choices). The utterance, although visual, can function as a directive illocutionary act because it is embedded in the context of anti-smoking public campaigns and visually represents a cigarette as a bullet. This metaphor implies a communicative function—discouraging smoking behavior by equating it with an act of self-inflicted violence. From this function, the speaker’s intention can be inferred as emerging from the pragmatic force and evaluative stance of the act itself, as to prevent smoking behavior by highlighting its deadly consequences through the conceptual mapping of smoking is shooting oneself. As for the directive kind, the discouraging attitude against smoking places the metaphorical act in a position that can be perceived by an annotator as the highly assertive end of the directive spectrum, referring to a prohibitive force similar to a command. This is achieved not through explicit verbal instruction but via the emotional and conceptual weight of the metaphorical image. The perlocutionary effects that can be experienced by an annotator can include emotions such as concern, discomfort, and shock as viewers process the visual equation of cigarettes with ammunition. The metaphor’s efficacy can be perceived as potentially high due to the stark, unambiguous nature of the weapon imagery and its cultural associations with death and violence. Visual clues reinforce the negative attitude through the absence of color, and the conceptual mapping between shooting a gun and smoking a cigarette is underlined by the presence of smoke. The overall visual tone of voice is dramatic and alarming, designed to interrupt habitual thinking about smoking through visceral impact rather than rational argument. This example of analysis is outlined in Figure 4.